# Large Knowledge Model: Perspectives and Challenges

### v1

@@ 3.5 Integration of Perception and Cognition
4 From LLM to LKM
Human knowledge, characterized by its complexity, is not optimally represented by mere sequences of free text. Our descriptions of the world often entail a variety of structured forms. The ultimate goal should to create more advanced Large Knowledge Model (LKM) proficient in handling and deciphering the myriad structures of knowledge representation. This section aims to illuminate a roadmap and identify the key challenges in developing such a Large Knowledge Model.
4.1 Decoupling World Knowledge from Language Models
4.2 Restructured Pretraining with Structure Knowledge
Numerous cognitive scientists posit that humans inherently favor structured thinking, tending to recall, reason, and plan via associative thinking. Recent research suggests that the way large models store knowledge lacks discernible patterns as like a KG StructGPT . In other words, those knowledge that should be structured correlated does not show the expected associations in the parameter space, making it difficult for the model to make feasible correlations among knowledge pieces, and generate the desired answers in a manner close to human cognition. There might be a sort of “cognitive gap” between the form of knowledge storage in large models and the organization of knowledge in the human mind. This cognitive gap becomes problematic in many critical scenarios when we expect the model to have more controllability, or to deal with deeper and more complex associative reasoning.
One way to alleviate this problem is to enhance the structure of the training corpus during the pretraining phase, known as restructure pretraining GraphGPT . That is, we can use automated or semi-automated methods to enhance the structured features of the text. For example, we can organize sentences more logically (essentially, CoT is a method to increase the structure and logic of texts). We can structure paragraphs more coherently, add links between words, create connections between training samples, or directly incorporating an external structured KG to structurally reinforce the entire training corpus.
By making the representation and organization of training data closer to the organization of knowledge in the human mind, we can reduce the cognitive gap between large models and human mind, thereby making the knowledge in large models more transparent, associative, and verifiable. Ultimately, this would paves a way to transform a LLM into a LKM, for which the structure of knowledge in a LKM aligns more naturally with human cognitive processes.
4.3 Large Commonsense Model
@@
Commonsense encompasses practical know-how and sound judgment regarding everyday matters, shared by nearly all humans. According to a definition from Marvin Minsky MinskyTheEmotionMachine , “commonsense knowledge includes facts about events occurring in time, about the effects of actions by the knower and others, about physical objects and how they are perceived, and about their properties and their relations to one another”.
The development of commonsense knowledge bases has been a pivotal challenge throughout AI’s history. This journey began with the early Cyc project Cyc , extended through ConceptNet ConceptNet as advocated by Marvin Minsky, and progressed to contemporary knowledge graph initiatives like Yago Yago , DBPedia DBPedia , Wikidata Wikidata , BabelNet BabelNet , ATOMIC Atomic , and ASTER ZHANG2022103740 . The consistent goal has been to equip AI with a comprehensive commonsense knowledge base. However, these common sense knowledge bases generally face challenges such as insufficient knowledge coverage and weak generalization capabilities of reasoning machines.
Large language models, such as ChatGPT, have demonstrated their potential in harboring vast commonsense knowledge and performing commonsense reasoning. Their principal strength lies in THE extensive knowledge coverage. However, they also grapple with challenges such as hallucination and the unreliability of reasoning results, and they tend to support only basic commonsense reasoning. A promising research direction for the future could involve integrating traditional commonsense knowledge bases with large language models to create an expansive large commonsense knowledge model. Such a model could potentially harness the strengths of both approaches to overcome their individual limitations.
@@ 4.4 Five-“A” Principle of LKM
In this final section, we strive to provide a general description on the concept of a Large Knowledge Model (LKM) and outline its pivotal characteristics. We outline the essential aspects of an LKM using a five-“A” framework as depicted in Figure. 2.
@@
  • Aligend with Knowledge: Finally, the entire system should be in close alignment with human values and ethics, which have evolved as a form of advanced knowledge accumulated throughout the history of humanity. This could be probably achieved through the creation of sophisticated ethics knowledge bases and knowledge alignment technologies zhang2023knowledgeable , which treat the ethic knowledge bases as aligning objectives.

### v2

@@ 4.4 Integration of Perception and Cognition
4.5 Commonsense Knowledge and Large World Model
A “world model” refers to a comprehensive model capable of accurately perceiving and understanding everything in the objective world and their complex relationships world-model . Unlike language models that are primarily used for understanding human language, and vision models that are mainly for understanding visual data, world models need to process data across various modalities in both temporal and spacial dimensions. They aim to form a cognition of the world and interact with the physical environment to perform specific tasks. World models are essential for achieving general AI. They must not only handle multimodal sensory data but also learn the spatiotemporal representations of the world effectively. Importantly, world models ultimately need to establish a common-sense knowledge model about the world, enabling them to accurately grasp and understand the complex relationships between all things in the world, so that the AI can respond and act correctly in the physical world.
@@ 4.5 Commonsense Knowledge and Large World Model
Knowledge models, especially common-sense knowledge models, are a crucial component of building world models. Commonsense knowledge encapsulates practical know-how and sound judgment about everyday situations, which is almost universally shared among humans. As defined by Marvin Minsky MinskyTheEmotionMachine , commonsense knowledge encompasses "facts about events occurring in time, about the effects of actions by the knower and others, about physical objects and how they are perceived, and about their properties and their relations to one another."
Historically, the creation of commonsense knowledge bases has been a key challenge in AI, beginning with early projects like Cyc Cyc and extending through initiatives such as ConceptNet ConceptNet , Yago Yago , DBPedia DBPedia , Wikidata Wikidata , BabelNet BabelNet , ATOMIC Atomic , and ASTER ZHANG2022103740 . These efforts aim to equip AI with robust commonsense knowledge bases. Yet, these systems often struggle with limited knowledge coverage and the reasoning capabilities of machines. Recent advancements in large language models like ChatGPT highlight significant potential in accumulating and applying commonsense knowledge, particularly due to their extensive knowledge coverage. Despite their strengths, these models face issues such as generating unreliable information (hallucination) and only supporting elementary commonsense reasoning tasks.
A promising direction for future research is to merge traditional commonsense knowledge bases with large language models to formulate a broad and integrated commonsense knowledge model. Such an approach could leverage the strengths of both systems to address their individual shortcomings, potentially leading to more robust and reliable Large World Model for AI in understanding and interacting with the real physical world.
@@ 4.6 Five-“A” Principle of LKM
In this final section, we strive to provide a general description on the concept of a Large Knowledge Model (LKM) and outline its pivotal characteristics. We outline the essential aspects of an LKM using a five-“A” framework as depicted in Figure 13.
@@
  • Aligend with Knowledge: Finally, the entire system should be in close alignment with human values and ethics, which have evolved as a form of advanced knowledge accumulated throughout the history of humanity. This could be probably achieved through the creation of sophisticated ethics knowledge bases and knowledge alignment technologies KnowPAT , which treat the ethic knowledge bases as aligning objectives.